My algebra is very, very rusty and I would like help with the problem below:
I have a 50% solution that I would like to mix with a 0.20 ml solution and achieve a final concentration of 6%. Please help me calculate the volume of the 50% solution that I will need to add to the 0.20ml to achieve a final concentration of 6% using an algebraic equation.

Answer by stanbon(48535)

My algebra is very, very rusty and I would like help with the problem below:
I have a 50% solution that I would like to mix with a 0.20 ml solution and achieve a final concentration of 6%. Please help me calculate the volume of the 50% solution that I will need to add to the 0.20ml to achieve a final concentration of 6% using an algebraic equation
Let amount of 50% solution be "x".
Active ingrediant in that solution is 0.5x ml
Equation:
active ingre. + active ingre. = active ingre.
0.5x + .20 = 0.6(x+0.2)
0.5x + 0.2 = 0.6x+0.12
0.1x = 0.2-0.12
0.1x = 0.08
x=0.8 ml
Amount of 50% solution you need is 0.8 ml
Cheers,
Stan H>